ZEEMAN PATTERNS IN FOURIER TRANSFORM RESOLVED FLUORESCENCE SPECTRA OF NiH

\underline{AMANDA J. ROSS}, PATRICK CROZET, HEATHER HARKER and CYRIL RICHARD, LASIM, Université Lyon 1 & CNRS, 43 Bd du 11 novembre 1918, F-69622 Villeurbanne, France ; and STEPHEN H. ASHWORTH, School of Chemistry, University of East Anglia, Norwich NR4 7TJ,UK.

We have found that dispersed laser induced fluorescence spectra of NiH produced in a discharge source are characterized by remarkably strong emission bands from upper state levels populated by collisional energy transfer.^{a} When the molecules are formed in the presence of a static magnetic field, magnetic response can be measured simultaneously for many transitions, under identical experimental conditions. Working with magnetic fields of 5000 - 10000 Gauss, the Zeeman patterns of many lines are at least partially resolved, at Doppler limited resolution. If the initial excitation process is chosen to be isotopically selective, the entire spectrum features emission from a single isotopologue, which simplifies analysis considerably. M_{J} selectivity is also possible in some cases. Measurements have been made so far for ^{58}NiH and 60NiH. The magnetic response of NiH in the low-lying states W_{1} ^{2}Π_{3/2}, X_{2} ^{2}Δ_{3/2}, and X_{1} ^{2}Δ_{5/2} states evidently changes with v′′. This illustrates the effects of mixing between states, described by the supermultiplet 3d^{9} complex model given by Gray et al.^{b} The intensity distributions of the M_{J} components in these spectra will be discussed.

aR. Vallon et al. Astrophys. J. \underline{696} 172 (2009)

bJ. A. Gray, M. Li, T. Nelis, and R. W. Field, J. Chem. Phys., \underline{95} 7164 (1991)